26 March 2011 LBC: continued testing with BP10 and checkout for science run

~1:30 John editing LBC header templates to match the latest master-slave ones. CRPIX1,2 values were not correct due to rotator center update on 20110129 which was not propagated to binocular SW.

~1:30 Notice OBJNAME missing from headers of 24 March.

1:45 Slew to flat field position to repeat flats of 24March. Trying to tweak relative exp time scalings: Y: +0.8 to +0.6 and I: -0.03 to -0.01 (good) Y counts seem to drop dramatically - is this right?

2:04 VI flats: 020516, 020544 B,R Flats have lots of stars in them so slewing to 12h 27deg and taking 30sec flats. Note that now after stopping an OB which was run simply to do the slew, the observer has the ability to play the next OB before the slew completes. We have to watch out for this.

2:20 BR flats: 022634,022631 but these have low counts even with T=30sec: 9000/5000.

2:32 Done with flats Now slewing back to 06:30 27 deg field which was rich with stars. Going to do initial FPIA, lbcrangebal, rotator center check, tilt check.

Note that every time we play and immediately stop an OB an error box appears which says something about status STOP and filter hardware error. This has been happening for months, but should be reported and the noted in the troubleshooting. It is harmless - click OK and go on - but disconcerting.

2:36 Starting FPIA. B 024203 has a poorly fit pupil to the bottom right. Converged in 5 iterations.

2:47 Slewing to WT10_301. FPIA converged in 3 iterations... Blue had too large Z6. Need to work on Z5/Z6. 25533/25531 are B/R pointing stars. 025922/025917 are B/R pointing stars after rangebal. The star images are on the rotator centers, so lbcrangebal is working after that change to the call to ptmodify.

3:00 Back to field at 06:36 and 27deg for a check of the tilt. Running dofpia.

3:10 rV superfoc at PA=0, then at PA=180 and then after dofpia (temperature changing rapidly) at PA=0 again. Analyzed focr1, first set, and see no evidence of tilt. Second set - seeing was much worse, but also no evidence of tilt.

3:48 dofpia and then SY superfoc. (Edited some Blue Superfoc for change to filter focus offsets but not all as yet!!!) Y best focus is -0.05 but we should check this when conditions stabilize. In conf file, Y best focus was set to -0.026. Changed all Red focus offsets by 0.01mm based on results for r-SLOAN on 9 March 2011 and tonight (but tonight is not yet stable... should monitor) _Note added later: I realized that I added, not subtracted, 0.01 for I-BESSEL, so 20110326 I images have focus offset -0.037 not -0.057. But I-B images looked round - leave as is, -0.037. (26 March afternoon)._

ROT CENTER B/R 0 deg: 035647/35648: B/R 180deg: 040021/040017. OLD ROT CEN = (1035, 2924) Blue and (1078, 2913) Red. Measured (1035, 2920) Blue and (1080, 2908.5) Red.

FPIA change John changed AstigScale from 220 to 200 in lbcfpia.cfg to reduce oscillations induced by seeing. Andrew & John discussing focus jump that occurs when seeing is too bad for spherical measurements (probably a small code change in LBCFPIA is needed).

ROT CENTER again at 0: B/R 044929/044926 and at 180: B/R 045146/045142 ===> I measure (1034, 2920) Blue and (1078.7, 2908.6) Red Decide on (1035, 2920) Blue and (1079, 2909) Red --- a small change --- about (0, -4) for both Blue and Red

05:00 Slewing to higher field - of galaxy NGC2841. ngc2841A1edit.ob where I edited the positions to try to put stars on both tech chips.

05:05 But first to pointing star ACT0361. Two exposures before dofpia. 050733/050730. Still see exposure time problems for Red Look at images: 025531R, 025917R. Compare B and R exposure times written to header for this pair (050733B, 050730R).

05:11 Blue pupils not well fit: 051142. Andrew & John sending some aberrations & watching dofpia.

05:22 052243/052240 are B/R on ACT0361. Second pair to confirm that lbcrangebal works.

05:26 ROTCen check here 0: 52807/52804 are B/R and then 180: 53002/52956.

EXPOSURE TIMES: NOTE that for the Rotator center OBs which do not send a telescope preset --- (-90 dec), the exposure times in headers for the Red/Blue sides are identical at 5.2sec. But for the fastextra observations, the exptime in header for Blue is 16.27 and for Red is 17.54 - these also do not send a preset. And for the pointing stars which do send a preset and are each 1sec, we see 1.27 for B and 2.54 for Red. Non capito!

5:40 Note that we want to have a look at the focus jumps that occurred when seeing got so bad that FPIA was no longer sending Z11 or Z22. See for example: 41616 and 41744 for red and 41923 and 42053 for blue.

05:44 Measure rotator centers for this field of ACT0361: Get (1032, 2918) Blue and (1079, 2906) Red. So both Red and Blue centers dropped by about 2 pixels in Y.

05:50-06:30 Observed galaxy NGC2841. A little too far E ... could have centered better (new OB ngc2841A1edit2_sci) has what I think is a better position. Guiding and focus corrections were sent for both B and R... not too many focus corrections for B, but I did see one. For R there were many. Observed 3 dither positions. I checked image quality with iqplot script for an I-BESSEL, R-BESSEL, z-SLOAN at R and U-BESSEL, V and B on B. All images had average ellip < 0.05 (0.036-0.045) and FWHM ~ 3.5 at I,Z to 5.5 at U.

06:39 Slewing to SN09nr field. Run FPIA on pointing star field: WT10_243. Then run lbcrangebal. OK.

06:45 Slew to SN09r and run FPIA... CRASHED because B image size is too small?! B image before that (065726) has bias step.
-rw-r--r-T 1 1127 1005  6508800 Mar 26 06:58 /newdata/lbcb.20110326.065726.fits
-rw-r--r-T 1 1127 1005    16384 Mar 26 06:59 /newdata/lbcb.20110326.065851.fits

Processing: /newdata/lbcb.20110326.065851.fits
EOF encountered reading extension header
Only 0 extensions processed
% GETIMAGE: Program caused arithmetic error: Floating illegal operand
% Execution halted at: GETIMAGE           72
  /home/lbcobs/LBCFPIA/lbcfpia/src/getimage.pro
%                      LBCFPIA           215
  /home/lbcobs/LBCFPIA/lbcfpia/src/lbcfpia.pro
%                      DOFPIA            444
  /home/lbcobs/LBCFPIA/lbcfpia/src/dofpia.pro
%                      $MAIN$          
% Program caused arithmetic error: Floating underflow
% Program caused arithmetic error: Floating overflow
% Program caused arithmetic error: Floating illegal operand

Jumps in Star Images on 074359 R Why?

07:47 continue to test guiding... There are stars in some positions on tech chips to enable focus corrections:

2011/03/26 08:01:35.426875 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 08:04:34.532043 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 08:04:38.702042 S R TRACKERS          Active Optics P03 +423.00000000 +0.00000000
2011/03/26 08:07:43.599062 S R TRACKERS          Active Optics P03 -18.00000000 +0.00000000
2011/03/26 08:10:38.383123 S B TRACKERS          Active Optics P03 +105.00000000 +0.00000000
2011/03/26 08:10:50.068145 S R TRACKERS          Active Optics P03 +58.00000000 +0.00000000
2011/03/26 08:16:55.846409 S R TRACKERS          Active Optics P03 -457.00000000 +0.00000000

08:37 GC in Canes Venatici. Going to this to watch tech chip focussing. SY, Vi, gr. 3 positions, 60sec each for a total of about 10 minutes for the OB. But first rangebal. Slew to ACT0294 and run dofpia. lbcrangebal: B/R 085140/085137 both IE/CA changed by about 11" and are close to what they were at the start of the night. ACT0294 again: 085426/085424.

08:58 At GC in CanVen, dofpia. Elevation = 85 deg now. Converged in 3 iterations.

09:04 Starting OB on GC in CanVen. 3 dither positions with 3 filters at each. Not so many focus corrections:
2011/03/26 09:05:55.508881 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:06:01.851053 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:07:34.658934 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:07:42.882699 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:09:12.724233 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:09:22.272825 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000

2011/03/26 09:10:58.453613 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:11:02.928815 S B TRACKERS          Active  Optics P03 +0.00000000 +0.00000000
2011/03/26 09:12:34.690466 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:12:43.249569 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:14:06.945322 S R TRACKERS          Active Optics P03 -360.00000000 +0.00000000
2011/03/26 09:14:22.540559 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000

2011/03/26 09:17:36.222653 S R TRACKERS          Active Optics P03 +486.00000000 +0.00000000
2011/03/26 09:17:43.857387 S B TRACKERS          Active Optics P03 -288.00000000 +0.00000000
2011/03/26 09:19:16.323276 S R TRACKERS          Active Optics P03 -543.00000000 +0.00000000
2011/03/26 09:19:24.129146 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000

09:21 Run dofpia again (20 minutes later) to see how well focus held up:

R 105 18 118 -96 147 -28 -19 0.90 0.13

B 796 -294 81 -3 -208 -89 -30 1.05 0.35

09:29 dofpia converged in 5 iterations. Running GC CanVen again, this time with an offset to the same position (3rd dither of previous OB) which did put stars on the focus chip last time.
2011/03/26 09:33:21.390293 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:33:32.418813 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:34:54.129737 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:35:08.912943 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:36:29.038649 S R TRACKERS          Active Optics P03 -804.00000000 +0.00000000
2011/03/26 09:36:48.915364 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:39:25.485106 S R TRACKERS          Active Optics P03 +337.00000000 +0.00000000
2011/03/26 09:39:25.875046 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:41:04.259253 S R TRACKERS          Active Optics P03 +94.00000000 +0.00000000
2011/03/26 09:42:41.669746 S R TRACKERS          Active Optics P03 -931.00000000 +0.00000000
2011/03/26 09:45:23.731910 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 09:45:25.001463 S R TRACKERS          Active Optics P03 +718.00000000 +0.00000000

09:44 Back to SN09 field while I prepare an OB with stars on both tech chips at all positions. WT10_294 for rangebal. FPIA --- took a while to converge. Observed pointing star. 100301/xxxx. lbcrangebal. Repeat pointing star observation. 100530/100527
 -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -479    129     29    -43    -15    -86    -17    y  0.69  0.22                      /newdata/lbcr.20110326.095312.fits  r-SLOAN 
   598     85    131   -179     45   -244    -28    y  0.78  0.31                      /newdata/lbcr.20110326.095453.fits  r-SLOAN 
  -614   -130    103     41   -161     51      5    y  0.92  0.27                      /newdata/lbcr.20110326.095635.fits  r-SLOAN 
  -633   -147   -110     -5    -17    165      0    y  0.83  0.29                      /newdata/lbcr.20110326.095814.fits  r-SLOAN 
   191   -184   -118   -121      2     92     18    y  0.92  0.15                      /newdata/lbcr.20110326.095955.fits  r-SLOAN 
   693      0      0      0      0   -396      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 
 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1050     11   -400    -86   -312   -127    -56    y  1.03  0.47                      /newdata/lbcb.20110326.095315.fits  V-BESSEL
  -123    228    186   -214    132     52      5    y  1.03  0.17                      /newdata/lbcb.20110326.095456.fits  V-BESSEL
   -43   -305    243    276    101    -74      8    y  1.05  0.20                      /newdata/lbcb.20110326.095638.fits  V-BESSEL
  -744   -198   -354   -225     -9    243     16    y  1.12  0.37                      /newdata/lbcb.20110326.095816.fits  V-BESSEL
   920    433    205    143   -210   -187    -32    y  1.15  0.43                      /newdata/lbcb.20110326.095958.fits  V-BESSEL
   878      0      0      0      0   -502      0    y  0.00  0.00                     dofpia_sph3_adjustment             

10:09 dofpia on sn09 field - now at elev=56deg. Still FPIA is oscillating now in focus:
 -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
    -2     -5    177   -175    118    -47      0    y  0.86  0.14                      /newdata/lbcr.20110326.100935.fits  r-SLOAN 
  -433     39   -190    153   -267     75      3    y  1.00  0.24                      /newdata/lbcr.20110326.101103.fits  r-SLOAN 
  -184   -146   -146   -137     16    -32      9    y  0.88  0.14                      /newdata/lbcr.20110326.101244.fits  r-SLOAN 
   669      0      0      0      0   -382      0    y  0.00  0.00                     dofpia_sph3_adjustment             

 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   718   -325    -68   -283     37    -23     -9    y  1.09  0.34                      /newdata/lbcb.20110326.100927.fits  V-BESSEL
  -700     26   -314    135     71    178     43    y  1.12  0.33                      /newdata/lbcb.20110326.101107.fits  V-BESSEL
   298    126    188    -59     70    -68    -23    y  1.13  0.17                      /newdata/lbcb.20110326.101246.fits  V-BESSEL
   865      0      0      0      0   -494      0    y  0.00  0.00                     dofpia_sph3_adjustment             

exited with "S" when Blue just missed converging for seeing 1.13" (z4=298). 

10:15 Start SN09 OB now at am = 1.217 and elev=55 deg. Seeing is good now! Sub-arcsecond. Will go on to do some standards and check filter focus offsets after this...

10:43 V-BESSEL 103515 Blue image has a blip to the NE of the star. This is visible for all bright stars. Telescope or mirror jump?? See attachment.
JMH: The brightest object on lbcbtec.20110326.103606_1 is a big (20") spiral galaxy with a bright nucleus. I suspect that seeing variations caused the guiding algorithm to get confused on this galaxy.

10:45 Done: below are focus adjustments and subsequent FPIA results on same field (first blue z4 was large, -1087nm; first red z4 was ~100nm)
2011/03/26 10:19:22.671283 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 10:22:19.999282 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2011/03/26 10:22:26.027909 S R TRACKERS          Active Optics P03 -47.00000000 +0.00000000
2011/03/26 10:25:31.430747 S R TRACKERS          Active Optics P03 +37.00000000 +0.00000000
2011/03/26 10:28:25.555271 S B TRACKERS          Active Optics P03 -186.00000000 +0.00000000
2011/03/26 10:28:40.200631 S R TRACKERS          Active Optics P03 -240.00000000 +0.00000000

  -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   180     34    385   -225    174    -52     -4    y  0.92  0.20                      /newdata/lbcr.20110326.105144.fits  r-SLOAN 
  -233     85    168    -53    -13     52     22    y  0.89  0.14                      /newdata/lbcr.20110326.105327.fits  r-SLOAN 
   290   -115    125     31    -54    -70      4    y  0.74  0.16                      /newdata/lbcr.20110326.105510.fits  r-SLOAN 
  -464   -184   -140   -242      3     73     -4    y  0.75  0.24                      /newdata/lbcr.20110326.105650.fits  r-SLOAN 
   -23    104   -163    338      3     52      9    y  0.95  0.18                      /newdata/lbcr.20110326.105834.fits  r-SLOAN 
   711      0      0      0      0   -406      0    y  0.00  0.00                     dofpia_sph3_adjustment             

 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1087     86   -240   -249   -102    -65    -10    y  1.06  0.46                      /newdata/lbcb.20110326.105145.fits  V-BESSEL
  -412     76   -203   -106    142     15      7    y  1.24  0.21                      /newdata/lbcb.20110326.105330.fits  V-BESSEL
   604      3    268     14     30    -49     -6    y  1.06  0.27                      /newdata/lbcb.20110326.105513.fits  V-BESSEL
  -658   -149    147   -110    -65     82     32    y  1.23  0.29                      /newdata/lbcb.20110326.105653.fits  V-BESSEL
   170    174   -267    121     71    -11     -7    y  1.22  0.15                      /newdata/lbcb.20110326.105837.fits  V-BESSEL
   922      0      0      0      0   -526      0    y  0.00  0.00                     dofpia_sph3_adjustment             

11:04 Slew to SA104 to observe in all filters. Looked photometric at sunset. Airmass of this standard is 1.8, elev=33deg. (Need more prepared Standard Star OBs). DOFPIA but stopped with "S" when close, given high airmass.

11:08 Running SA104 OB. 110855/110845.... Red channel execution error. Went through all filters at least once,so will move on to SA107 after powering down/up the cameras.

11:33 Slew to SA107 and start DOFPIA. Airmass = 1.221, elev=55 deg. Exit with "S". More exposure time weirdness. Red r-SLOAN,Y, should be 3.0sec, but header reads 4.5sec and Blue U,S should be 3.0sec and header reads (as expected) 3.2sec.

12:00 FPIA but seeing estimate is large in blue especially (1.6-1.8). Both Blue and Red pupils have a gradient from top (bright) to bottom (dimmer). Swing arm opening toward about 5 oclock. Moon-object angle is 46 deg --- could this be a problem??

12:09 Repeating SA107 OB. Now at Airmass 1.28 and elevation 51 deg. Note we are in astronomical twilight.

12:16 Slew to a field further from moon to try focus offsets... I am not so sure about the seeing estimates from the previous FPIA when the pupils were not uniformly illuminated. RA=18, DEC=45. (see attachment).

12:28 FPIA on this field... these pupils look very detailed - seeing is good. FPIA converged in 4 iterations. Superfoc_ig (focus center was off), Superfoc_rV, Superfoc_ig (with edited OB). Getting too light. The g focus looks OK, but it is not a good fit. Need to look at sequence more closely and also at science images taken through these filters.

12:46 Flat fields at V and I. Counts start around 13.5k, 16k in V, I. Next a few flats at Sz. Last two at S saturated.

12:55 Closing --- while closing, taking some biases. *In evening, Red biases taken with light (dome open) showed about 2x as many counts as the bias level and a strong gradient across short dimension of chip. But now, with dome still open, these Red biases look fine - about 340 counts and flat, which is normal. Blue bias has some banding (but this image was taken while slewing/closing up).

13:01 Turning LBC off. Summary
  • One RPC error on the Red Camera last night. Two truncated Blue images (lbcb.20110326.081123.fits and lbcb.20110326.102258.fits --- both have sizes of 16384 bytes instead of 84983040 bytes). Otherwise, LBC worked well.
  • At least one image on which we saw a jump (perhaps I spotted another one earlier on). But image quality looks good on the whole.
  • Some variability in estimated seeing from FPIA - I think this may be real, but the pupil shapes and illuminations are sometimes weird.
  • Unfortunately, no filter focus offsets due to the perceived seeing variations - at the end of the night tried a few focus sequences, but it was getting light and the fits were not good. Data taken through many filters, so we do have images on which to check image shapes - from what I could tell, these looked good (but is g-SLOAN focus offset good?)
  • Observed some standard stars SA104 at low elevation and SA107 at higher elevation.
  • Exposure time written to red headers is longer than what is requested by 1.3 seconds. But I did not see this in all instances...(not on focus sequence but on fastextra, standards, flats).
  • Running "allseeing" on images (set time cut to 1sec to get the standards even though speckle will dominate). See attached output files (two for blue).
  • Observed NGC2841, Globular Cluster in Canes Venatici, SN09nr, SA104, SA107.

-- OlgaKuhn - 26 Mar 2011
  • Error box that appears when an OB is played & stopped:
    OBSTOP_error.jpg

  • Blue Image --- small blip to NE of bright stars. What happened?:
    Blue_103515_tel_or_mirr_jump.jpg

  • Illumination gradient over pupils: est seeing 1.1" R and 1.8" B:
    illumgradient.jpg

  • Pupils in good seeing (0.76/0.8" for B/R) Odd extension on both at about 7 o'clock.:
    goodseeing.jpg

  • 20110326.fwhm_Blue2: allseeing run on second part of Blue images (truncated images crashed script)

I Attachment Action Size Date Who Comment
20110326.fwhm_Blue2fwhm_Blue2 20110326.fwhm_Blue2 manage 5 K 27 Mar 2011 - 06:20 OlgaKuhn allseeing run on second part of Blue images (truncated images crashed script)
20110326_Blue1.fwhmfwhm 20110326_Blue1.fwhm manage 2 K 27 Mar 2011 - 06:19 OlgaKuhn allseeing run on first part of Blue images
20110326_Red.fhwmfhwm 20110326_Red.fhwm manage 10 K 27 Mar 2011 - 06:20 OlgaKuhn allseeing run on Red images
Blue_103515_tel_or_mirr_jump.jpgjpg Blue_103515_tel_or_mirr_jump.jpg manage 93 K 26 Mar 2011 - 10:47 OlgaKuhn Blue Image --- small blip to NE of bright stars. What happened?
OBSTOP_error.jpgjpg OBSTOP_error.jpg manage 11 K 26 Mar 2011 - 08:21 OlgaKuhn Error box that appears when an OB is played & stopped
goodseeing.jpgjpg goodseeing.jpg manage 32 K 26 Mar 2011 - 13:33 OlgaKuhn Pupils in good seeing (0.76/0.8" for B/R 122638/122635) Odd extension on both at about 7 o'clock.
illumgradient.jpgjpg illumgradient.jpg manage 17 K 26 Mar 2011 - 12:24 OlgaKuhn Illumination gradient over pupils: est seeing 1.1" R (120100) and 1.8" B (120103)
Topic revision: r9 - 27 Mar 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback